# IR2: Information Regularization for Information Retrieval thanks:   Accepted by LREC-COLING 2024

### v1

@@ Abstract
Effective information retrieval (IR) in settings with limited training data, particularly for complex queries, remains a challenging task. This paper introduces IR2, Information Regularization for Information Retrieval, a technique for reducing overfitting during synthetic data generation. This approach, representing a novel application of regularization techniques in synthetic data creation for IR, is tested on three recent IR tasks characterized by complex queries: DORIS-MAE, ArguAna, and WhatsThatBook. Experimental results indicate that our regularization techniques not only outperform previous synthetic query generation methods on the tasks considered but also reduce cost by up to 50%. Furthermore, this paper categorizes and explores three regularization methods at different stages of the query synthesis pipeline—input, prompt, and output—each offering varying degrees of performance improvement compared to models where no regularization is applied. This provides a systematic approach for optimizing synthetic data generation in data-limited, complex-query IR scenarios. All code, prompts and synthetic data are available at https://github.com/Info-Regularization/Information-Regularization.
Keywords: Information Retrieval, Data Augmentation, Synthetic Query Generation, Regularization, Contrastive Learning, Large Language Models
[…]
@@
Jianyou Wangthanks:   Equal contribution  Kaicheng Wang22footnotemark: 2  Xiaoyue Wang22footnotemark: 2 Weili Cao, Ramamohan Paturithanks:   Equal contribution  Leon Bergen33footnotemark: 3
Laboratory for Emerging Intelligence
University of California, San Diego
{ jiw101, kaw036, xiw027, w2cao, rpaturi, lbergen }@ucsd.edu
Abstract content
1.   Introduction
@@
Figure 1: Performance of synthetic data generation methods on complex IR benchmarks. The $\triangle$, $\square$ Gao et al. (2021), and ◯ Dai et al. (2022) icons represent baselines. The other four icons Refer to caption Refer to caption Refer to caption Refer to caption denote IR2 approaches, indicating the performance of models after fine-tuning on information-regularized synthetic datasets. Metrics are chosen based on standard practice for the three benchmarks. Model performances are averaged across all models used in experiments.
@@ 2.   Related Work
IR tasks have typically centered on retrieving relevant passages to answer simple, sentence-level queries, as seen in foundational datasets like MS MARCO Nguyen et al. (2016) and NQ (Kwiatkowski et al., 2019). The BEIR dataset Thakur et al. (2021) combines 19 IR tasks, most of which are sentence-level, including specific challenges like TREC-COVID (Voorhees et al., 2020) and SCIFACT Wadden et al. (2020). Several recent datasets contain complex, paragraph-length queries. DORIS-MAE (Wang et al., 2023) poses the task of retrieving scientific abstracts given complex scientific research questions. In the WhatsThatBook dataset Lin et al. (2023), tip-of-the-tongue user queries are used to retrieve book descriptions. In ArguAna (Wachsmuth et al., 2018), paragraph-length arguments are used to retrieve counterarguments.
@@ 3.   Methodology
3.1.   Synthetic Query Generation
@@
3.2.   Baseline Query Generation
@@ 3.3.   Information Regularization
3.3.1.   Input Document Regularization
@@ 3.3.2.   Instruction Regularization
3.3.3.   Output Query Regularization
@@ 4.   Experiments
4.1.   Benchmarks
WhatsThatBook (WTB) (Lin et al., 2023) contains 14,441 queries from users trying to recall specific books, paired with book titles and descriptions. The task is to match each of the 1,445 test set queries with the correct descriptions. We use descriptions from 4,000 unique books, exclusive from the test set, for synthetic query generation.
@@ 4.1.   Benchmarks
4.2.   Models
We experiment with four models: E5-Large-v2 Wang et al. (2022c) (355M), SimCSE-Large Gao et al. (2021) (355M), RoBERTa-Large Liu et al. (2019) (355M), and SPECTER-v2 Cohan et al. (2020) (110M). E5-Large-v2 (denoted as E5), a contrastively trained model, demonstrates performance on par with the state-of-the-art ada-002 Greene et al. (2022) on BEIR Thakur et al. (2021). For SimCSE-Large (denoted as SimCSE), which uses self-supervised training with dropout masks, we choose the top-performing checkpoint on the STS benchmark (Agirre et al., 2013). RoBERTa-Large is denoted as RoBERTa. SPECTER-v2 is specialized for scientific document representation. Embeddings are derived according to each model’s specifications, using pooling strategies, and retrieval and reranking are based on similarity measures (L2 distance for SPECTER-v2; cosine similarity for others).
4.3.   Baselines
@@
Method
R@5
R@20
NDCG@10
MRR@10
R@5
R@20
NDCG@10
MRR@10
Pretrained
18.41
27.06
15.25
13.07
2.84
5.33
2.17
1.73
Contrastive Fine-tuning
37.25
49.85
33.16
29.87
23.59
35.27
20.69
18.10
Promptagator
41.53
54.66
37.07
33.56
27.23
39.14
23.69
20.78
Doc40% reg
41.40
53.69 $\downarrow$ -0.97
37.01
33.72
28.30 $\uparrow$ 1.07
40.89 $\uparrow$ 1.75
24.79 $\uparrow$ 1.10
21.81 $\uparrow$ 1.03
Queryreg $\circ$ Doc40% reg
45.50 $\uparrow$ 3.97
58.08 $\uparrow$ 3.42
40.50 $\uparrow$ 3.43
37.01 $\uparrow$ 3.45
32.83 $\uparrow$ 5.60
46.38 $\uparrow$ 7.24
29.05 $\uparrow$ 5.36
25.82 $\uparrow$ 5.04
Doc60% reg
44.15 $\uparrow$ 2.62
55.69 $\uparrow$ 1.03
39.64 $\uparrow$ 2.57
36.16 $\uparrow$ 2.60
28.67 $\uparrow$ 1.44
41.49 $\uparrow$ 2.35
25.14 $\uparrow$ 1.45
22.06 $\uparrow$ 1.28
Queryreg $\circ$ Doc60% reg
46.23 $\uparrow$ 4.70
59.76 $\uparrow$ 5.10
41.58 $\uparrow$ 4.51
38.13 $\uparrow$ 4.57
34.30 $\uparrow$ 7.07
47.52 $\uparrow$ 8.38
30.15 $\uparrow$ 6.46
26.83 $\uparrow$ 6.05
Doc80% reg
28.30 $\downarrow$ -13.23
39.85 $\downarrow$ -14.81
24.05 $\downarrow$ -13.02
21.13 $\downarrow$ -12.43
21.74 $\downarrow$ -5.49
33.16 $\downarrow$ -5.98
18.91 $\downarrow$ -4.78
16.35 $\downarrow$ -4.43
Queryreg $\circ$ Doc80% reg
44.80 $\uparrow$ 3.27
58.21 $\uparrow$ 3.55
40.29 $\uparrow$ 3.22
36.94 $\uparrow$ 3.38
33.69 $\uparrow$ 6.46
46.23 $\uparrow$ 7.09
29.41 $\uparrow$ 5.72
26.20 $\uparrow$ 5.42
Instrreg
42.15 $\uparrow$ 0.62
53.02 $\downarrow$ -1.64
37.33
33.89 $\uparrow$ 0.33
28.12 $\uparrow$ 0.89
40.80 $\uparrow$ 1.66
24.51 $\uparrow$ 0.82
21.46 $\uparrow$ 0.68
Queryreg $\circ$ Instrreg
44.90 $\uparrow$ 3.37
56.28 $\uparrow$ 1.62
39.29 $\uparrow$ 2.22
35.79 $\uparrow$ 2.23
33.09 $\uparrow$ 5.86
46.09 $\uparrow$ 6.95
29.28 $\uparrow$ 5.59
26.06 $\uparrow$ 5.28
@@
Method
R@5
R@20
NDCG@10
MRR@10
R@5
R@20
NDCG@10
MRR@10
Pretrained
16.96
24.98
14.58
12.67
2.98
5.26
2.43
1.87
Contrastive Fine-tuning
26.78
37.15
23.11
20.42
7.85
12.56
6.36
5.35
Promptagator
30.97
44.26
27.15
23.99
6.56
11.81
5.72
4.88
Doc40% reg
32.10 $\uparrow$ 1.13
43.07 $\downarrow$ -1.19
28.30 $\uparrow$ 1.15
25.45 $\uparrow$ 1.46
9.98 $\uparrow$ 3.42
16.12 $\uparrow$ 4.31
8.55 $\uparrow$ 2.83
7.32 $\uparrow$ 2.44
Queryreg $\circ$ Doc40% reg
34.63 $\uparrow$ 3.66
47.44 $\uparrow$ 3.18
30.64 $\uparrow$ 3.49
27.51 $\uparrow$ 3.52
10.84 $\uparrow$ 4.28
16.26 $\uparrow$ 4.45
8.81 $\uparrow$ 3.09
7.57 $\uparrow$ 2.69
Doc60% reg
33.03 $\uparrow$ 2.06
43.96
28.87 $\uparrow$ 1.72
25.97 $\uparrow$ 1.98
8.64 $\uparrow$ 2.08
14.49 $\uparrow$ 2.68
7.46 $\uparrow$ 1.74
6.26 $\uparrow$ 1.38
Queryreg $\circ$ Doc60% reg
34.84 $\uparrow$ 3.87
47.08 $\uparrow$ 2.82
30.77 $\uparrow$ 3.62
27.64 $\uparrow$ 3.65
7.54 $\uparrow$ 0.98
13.29 $\uparrow$ 1.48
6.55 $\uparrow$ 0.83
5.48 $\uparrow$ 0.60
Doc80% reg
19.75 $\downarrow$ -11.22
29.89 $\downarrow$ -14.37
16.73 $\downarrow$ -10.42
14.33 $\downarrow$ -9.66
7.39 $\uparrow$ 0.83
13.19 $\uparrow$ 1.38
6.46 $\uparrow$ 0.74
5.31 $\uparrow$ 0.43
Queryreg $\circ$ Doc80% reg
33.76 $\uparrow$ 2.79
45.63 $\uparrow$ 1.37
29.34 $\uparrow$ 2.19
26.26 $\uparrow$ 2.27
8.14 $\uparrow$ 1.58
14.69 $\uparrow$ 2.88
7.33 $\uparrow$ 1.61
6.13 $\uparrow$ 1.25
Instrreg
33.81 $\uparrow$ 2.84
45.79 $\uparrow$ 1.53
29.74 $\uparrow$ 2.59
26.79 $\uparrow$ 2.80
10.49 $\uparrow$ 3.93
16.80 $\uparrow$ 4.99
9.30 $\uparrow$ 3.58
8.07 $\uparrow$ 3.19
Queryreg $\circ$ Instrreg
37.34 $\uparrow$ 6.37
50.45 $\uparrow$ 6.19
33.18 $\uparrow$ 6.03
30.02 $\uparrow$ 6.03
11.08 $\uparrow$ 4.52
17.36 $\uparrow$ 5.55
9.59 $\uparrow$ 3.87
8.28 $\uparrow$ 3.40
@@ 4.5.   Training Methods
5.   Results
5.1.   Results on WhatsThatBook

### v2

@@ Abstract
Effective information retrieval (IR) in settings with limited training data, particularly for complex queries, remains a challenging task. This paper introduces IR2, Information Regularization for Information Retrieval, a technique for reducing overfitting during synthetic data generation. This approach, representing a novel application of regularization techniques in synthetic data creation for IR, is tested on three recent IR tasks characterized by complex queries: DORIS-MAE, ArguAna, and WhatsThatBook. Experimental results indicate that our regularization techniques not only outperform previous synthetic query generation methods on the tasks considered but also reduce cost by up to 50%. Furthermore, this paper categorizes and explores three regularization methods at different stages of the query synthesis pipeline—input, prompt, and output—each offering varying degrees of performance improvement compared to models where no regularization is applied. This provides a systematic approach for optimizing synthetic data generation in data-limited, complex-query IR scenarios. All code, prompts and synthetic data are available at https://github.com/kwang927/Information-Regularization.
@@
Jianyou Wang thanks: Equal contribution  Kaicheng Wang22footnotemark: 2  Xiaoyue Wang22footnotemark: 2  Weili Cao Ramamohan Paturi thanks: Equal Senior Authors  Leon Bergen33footnotemark: 3 Laboratory for Emerging Intelligence University of California, San Diego {jiw101, kaw036, xiw027, w2cao, rpaturi, lbergen}@ucsd.edu
1 Introduction
@@
Figure 1: Performance of synthetic data generation methods on complex IR benchmarks. The $\triangle$, $\square$ Gao et al. (2021), and ⃝ Dai et al. (2022) icons represent baselines. The other four icons Refer to caption Refer to caption Refer to caption Refer to caption denote IR2 approaches, indicating the performance of models after fine-tuning on information-regularized synthetic datasets. Metrics are chosen based on standard practice for the three benchmarks. Model performances are averaged across all models used in experiments.
@@ 2 Related Work
IR tasks have typically centered on retrieving relevant passages to answer simple, sentence-level queries, as seen in foundational datasets like MS MARCO Nguyen et al. (2016) and NQ Kwiatkowski et al. (2019). The BEIR dataset Thakur et al. (2021) combines 19 IR tasks, most of which are sentence-level, including specific challenges like TREC-COVID Voorhees et al. (2021) and SCIFACT Wadden et al. (2020). Several recent datasets contain complex, paragraph-length queries. DORIS-MAE Wang et al. (2023) poses the task of retrieving scientific abstracts given complex scientific research questions. In the WhatsThatBook dataset Lin et al. (2023), tip-of-the-tongue user queries are used to retrieve book descriptions. In ArguAna Wachsmuth et al. (2018), paragraph-length arguments are used to retrieve counterarguments.
@@ 3 Methodology
3.1 Synthetic Query Generation
@@
3.2 Baseline Query Generation
@@ 3.3 Information Regularization
3.3.1 Input Document Regularization
@@ 3.3.2 Instruction Regularization
3.3.3 Output Query Regularization
@@ 4 Experiments
4.1 Benchmarks
WhatsThatBook (WTB) Lin et al. (2023) contains 14,441 queries from users trying to recall specific books, paired with book titles and descriptions. The task is to match each of the 1,445 test set queries with the correct descriptions. We use descriptions from 4,000 unique books, exclusive from the test set, for synthetic query generation.
@@ 4.1 Benchmarks
4.2 Models
We experiment with four models: E5-Large-v2 Wang et al. (2022c) (355M), SimCSE-Large Gao et al. (2021) (355M), RoBERTa-Large Liu et al. (2019) (355M), and SPECTER-v2 Cohan et al. (2020) (110M). E5-Large-v2 (denoted as E5), a contrastively trained model, demonstrates performance on par with the state-of-the-art ada-002 Greene et al. (2022) on BEIR Thakur et al. (2021). For SimCSE-Large (denoted as SimCSE), which uses self-supervised training with dropout masks, we choose the top-performing checkpoint on the STS benchmark Agirre et al. (2013). RoBERTa-Large is denoted as RoBERTa. SPECTER-v2 is specialized for scientific document representation. Embeddings are derived according to each model’s specifications, using pooling strategies, and retrieval and reranking are based on similarity measures (L2 distance for SPECTER-v2; cosine similarity for others).
4.3 Baselines
@@
Method R@5 R@20 NDCG@10 MRR@10 R@5 R@20 NDCG@10 MRR@10
Pretrained 18.41 27.06 15.25 13.07 2.84 5.33 2.17 1.73
Contrastive Fine-tuning 37.25 49.85 33.16 29.87 23.59 35.27 20.69 18.10
Promptagator 41.53 54.66 37.07 33.56 27.23 39.14 23.69 20.78
Doc40% reg 41.40 53.69 $\downarrow$ -0.97 37.01 33.72 28.30 $\uparrow$ 1.07 40.89 $\uparrow$ 1.75 24.79 $\uparrow$ 1.10 21.81 $\uparrow$ 1.03
Queryreg $\circ$ Doc40% reg 45.50 $\uparrow$ 3.97 58.08 $\uparrow$ 3.42 40.50 $\uparrow$ 3.43 37.01 $\uparrow$ 3.45 32.83 $\uparrow$ 5.60 46.38 $\uparrow$ 7.24 29.05 $\uparrow$ 5.36 25.82 $\uparrow$ 5.04
Doc60% reg 44.15 $\uparrow$ 2.62 55.69 $\uparrow$ 1.03 39.64 $\uparrow$ 2.57 36.16 $\uparrow$ 2.60 28.67 $\uparrow$ 1.44 41.49 $\uparrow$ 2.35 25.14 $\uparrow$ 1.45 22.06 $\uparrow$ 1.28
Queryreg $\circ$ Doc60% reg 46.23 $\uparrow$ 4.70 59.76 $\uparrow$ 5.10 41.58 $\uparrow$ 4.51 38.13 $\uparrow$ 4.57 34.30 $\uparrow$ 7.07 47.52 $\uparrow$ 8.38 30.15 $\uparrow$ 6.46 26.83 $\uparrow$ 6.05
Doc80% reg 28.30 $\downarrow$ -13.23 39.85 $\downarrow$ -14.81 24.05 $\downarrow$ -13.02 21.13 $\downarrow$ -12.43 21.74 $\downarrow$ -5.49 33.16 $\downarrow$ -5.98 18.91 $\downarrow$ -4.78 16.35 $\downarrow$ -4.43
Queryreg $\circ$ Doc80% reg 44.80 $\uparrow$ 3.27 58.21 $\uparrow$ 3.55 40.29 $\uparrow$ 3.22 36.94 $\uparrow$ 3.38 33.69 $\uparrow$ 6.46 46.23 $\uparrow$ 7.09 29.41 $\uparrow$ 5.72 26.20 $\uparrow$ 5.42
Instrreg 42.15 $\uparrow$ 0.62 53.02 $\downarrow$ -1.64 37.33 33.89 $\uparrow$ 0.33 28.12 $\uparrow$ 0.89 40.80 $\uparrow$ 1.66 24.51 $\uparrow$ 0.82 21.46 $\uparrow$ 0.68
Queryreg $\circ$ Instrreg 44.90 $\uparrow$ 3.37 56.28 $\uparrow$ 1.62 39.29 $\uparrow$ 2.22 35.79 $\uparrow$ 2.23 33.09 $\uparrow$ 5.86 46.09 $\uparrow$ 6.95 29.28 $\uparrow$ 5.59 26.06 $\uparrow$ 5.28
@@
Method R@5 R@20 NDCG@10 MRR@10 R@5 R@20 NDCG@10 MRR@10
Pretrained 16.96 24.98 14.58 12.67 2.98 5.26 2.43 1.87
Contrastive Fine-tuning 26.78 37.15 23.11 20.42 7.85 12.56 6.36 5.35
Promptagator 30.97 44.26 27.15 23.99 6.56 11.81 5.72 4.88
Doc40% reg 32.10 $\uparrow$ 1.13 43.07 $\downarrow$ -1.19 28.30 $\uparrow$ 1.15 25.45 $\uparrow$ 1.46 9.98 $\uparrow$ 3.42 16.12 $\uparrow$ 4.31 8.55 $\uparrow$ 2.83 7.32 $\uparrow$ 2.44
Queryreg $\circ$ Doc40% reg 34.63 $\uparrow$ 3.66 47.44 $\uparrow$ 3.18 30.64 $\uparrow$ 3.49 27.51 $\uparrow$ 3.52 10.84 $\uparrow$ 4.28 16.26 $\uparrow$ 4.45 8.81 $\uparrow$ 3.09 7.57 $\uparrow$ 2.69
Doc60% reg 33.03 $\uparrow$ 2.06 43.96 28.87 $\uparrow$ 1.72 25.97 $\uparrow$ 1.98 8.64 $\uparrow$ 2.08 14.49 $\uparrow$ 2.68 7.46 $\uparrow$ 1.74 6.26 $\uparrow$ 1.38
Queryreg $\circ$ Doc60% reg 34.84 $\uparrow$ 3.87 47.08 $\uparrow$ 2.82 30.77 $\uparrow$ 3.62 27.64 $\uparrow$ 3.65 7.54 $\uparrow$ 0.98 13.29 $\uparrow$ 1.48 6.55 $\uparrow$ 0.83 5.48 $\uparrow$ 0.60
Doc80% reg 19.75 $\downarrow$ -11.22 29.89 $\downarrow$ -14.37 16.73 $\downarrow$ -10.42 14.33 $\downarrow$ -9.66 7.39 $\uparrow$ 0.83 13.19 $\uparrow$ 1.38 6.46 $\uparrow$ 0.74 5.31 $\uparrow$ 0.43
Queryreg $\circ$ Doc80% reg 33.76 $\uparrow$ 2.79 45.63 $\uparrow$ 1.37 29.34 $\uparrow$ 2.19 26.26 $\uparrow$ 2.27 8.14 $\uparrow$ 1.58 14.69 $\uparrow$ 2.88 7.33 $\uparrow$ 1.61 6.13 $\uparrow$ 1.25
Instrreg 33.81 $\uparrow$ 2.84 45.79 $\uparrow$ 1.53 29.74 $\uparrow$ 2.59 26.79 $\uparrow$ 2.80 10.49 $\uparrow$ 3.93 16.80 $\uparrow$ 4.99 9.30 $\uparrow$ 3.58 8.07 $\uparrow$ 3.19
Queryreg $\circ$ Instrreg 37.34 $\uparrow$ 6.37 50.45 $\uparrow$ 6.19 33.18 $\uparrow$ 6.03 30.02 $\uparrow$ 6.03 11.08 $\uparrow$ 4.52 17.36 $\uparrow$ 5.55 9.59 $\uparrow$ 3.87 8.28 $\uparrow$ 3.40
@@ 4.5 Training Methods
E5-Large-v2 RoBERTa
Method R@5 R@20 NDCG@10 MRR@10 R@5 R@20 NDCG@10 MRR@10
Pretrained 14.67 42.15 71.98 14.34 11.96 34.77 66.86 8.56
Contrastive Fine-tuning 15.28 39.87 72.51 13.43 14.75 43.32 73.55 20.79
Promptagator 16.18 45.59 73.95 14.83 13.03 42.88 72.23 16.82
Doc40% reg 14.90 $\downarrow$ -1.28 46.55 $\uparrow$ 0.96 74.28 $\uparrow$ 0.33 14.87 12.54 44.89 $\uparrow$ 2.01 72.94 $\uparrow$ 0.71 18.16 $\uparrow$ 1.34
Queryreg $\circ$ Doc40% reg 15.85 45.67 73.96 16.10 $\uparrow$ 1.27 14.66 $\uparrow$ 1.63 47.41 $\uparrow$ 4.53 74.04 $\uparrow$ 1.81 18.81 $\uparrow$ 1.99
Doc60% reg 15.23 $\downarrow$ -0.95 46.57 $\uparrow$ 0.98 74.98 $\uparrow$ 1.03 15.14 14.11 $\uparrow$ 1.08 43.99 $\uparrow$ 1.11 72.94 $\uparrow$ 0.71 16.84
Queryreg $\circ$ Doc60% reg 15.31 $\downarrow$ -0.87 48.16 $\uparrow$ 2.57 74.90 $\uparrow$ 0.95 15.43 $\uparrow$ 0.60 14.75 $\uparrow$ 1.72 47.25 $\uparrow$ 4.37 74.61 $\uparrow$ 2.38 20.48 $\uparrow$ 3.66
Doc80% reg 14.68 $\downarrow$ -1.50 43.89 $\downarrow$ -1.70 74.13 $\uparrow$ 0.18 13.51 $\downarrow$ -1.32 14.47 $\uparrow$ 1.44 42.47 72.61 $\uparrow$ 0.38 15.10 $\downarrow$ -1.72
Queryreg $\circ$ Doc80% reg 15.63 $\downarrow$ -0.55 46.75 $\uparrow$ 1.16 75.08 $\uparrow$ 1.13 16.62 $\uparrow$ 1.79 15.84 $\uparrow$ 2.81 46.96 $\uparrow$ 4.08 74.71 $\uparrow$ 2.48 20.49 $\uparrow$ 3.67
Instrreg 16.38 46.73 $\uparrow$ 1.14 74.42 $\uparrow$ 0.47 17.06 $\uparrow$ 2.23 13.80 $\uparrow$ 0.77 45.78 $\uparrow$ 2.90 73.53 $\uparrow$ 1.30 18.24 $\uparrow$ 1.42
Queryreg $\circ$ Instrreg 15.76 $\downarrow$ -0.42 48.46 $\uparrow$ 2.87 76.02 $\uparrow$ 2.07 20.86 $\uparrow$ 6.03 15.20 $\uparrow$ 2.17 47.44 $\uparrow$ 4.56 74.08 $\uparrow$ 1.85 19.31 $\uparrow$ 2.49
SimCSE SPECTER-v2
Method R@5 R@20 NDCG@10 MRR@10 R@5 R@20 NDCG@10 MRR@10
Pretrained 14.38 41.83 70.81 24.83 13.59 41.92 71.46 21.74
Contrastive Fine-tuning 15.60 44.99 73.18 22.41 14.56 45.64 71.46 20.49
Promptagator 16.35 46.69 73.97 16.54 14.13 39.22 71.55 22.37
Doc40% reg 17.70 $\uparrow$ 1.35 45.30 $\downarrow$ -1.39 74.49 $\uparrow$ 0.52 16.49 15.39 $\uparrow$ 1.26 43.39 $\uparrow$ 4.17 72.11 $\uparrow$ 0.56 21.66 $\downarrow$ -0.71
Queryreg $\circ$ Doc40% reg 16.06 45.46 $\downarrow$ -1.23 73.23 $\downarrow$ -0.74 16.56 14.19 42.66 $\uparrow$ 3.44 71.49 21.40 $\downarrow$ -0.97
Doc60% reg 17.08 $\uparrow$ 0.73 45.33 $\downarrow$ -1.36 74.70 $\uparrow$ 0.73 17.75 $\uparrow$ 1.21 15.41 $\uparrow$ 1.28 44.19 $\uparrow$ 4.97 72.81 $\uparrow$ 1.26 21.35 $\downarrow$ -1.02
Queryreg $\circ$ Doc60% reg 17.04 $\uparrow$ 0.69 46.68 74.65 $\uparrow$ 0.68 18.32 $\uparrow$ 1.78 14.95 $\uparrow$ 0.82 43.30 $\uparrow$ 4.08 72.19 $\uparrow$ 0.64 21.60 $\downarrow$ -0.77
Doc80% reg 15.01 $\downarrow$ -1.34 42.53 $\downarrow$ -4.16 73.69 $\downarrow$ -0.28 18.68 $\uparrow$ 2.14 17.02 $\uparrow$ 2.89 44.97 $\uparrow$ 5.75 73.94 $\uparrow$ 2.39 22.78 $\uparrow$ 0.41
Queryreg $\circ$ Doc80% reg 15.92 45.70 $\downarrow$ -0.99 73.86 18.54 $\uparrow$ 2.00 14.70 $\uparrow$ 0.57 44.66 $\uparrow$ 5.44 73.07 $\uparrow$ 1.52 22.65
Instrreg 16.50 46.68 75.60 $\uparrow$ 1.63 19.71 $\uparrow$ 3.17 14.73 $\uparrow$ 0.60 41.50 $\uparrow$ 2.28 73.06 $\uparrow$ 1.51 21.05 $\downarrow$ -1.32
Queryreg $\circ$ Instrreg 16.06 47.61 $\uparrow$ 0.92 75.43 $\uparrow$ 1.46 22.14 $\uparrow$ 5.60 14.50 $\uparrow$ 0.37 41.44 $\uparrow$ 2.22 72.46 $\uparrow$ 0.91 22.26
5 Results
5.1 Results on WhatsThatBook